OpenStudy (anonymous):

Are these lines parallel perpendicular or neither? 1. Y=-4x+3 -2x+8y=5 2. Y= -1/4x + 10 -2x + 8y = 6

OpenStudy (compassionate):

Hi, parallel lines will have the same slope. Perpendicular lines will have opposite slopes. y = -4x + 3 and -2x + 8y = 5 The first one is already solved for y, so just solve the second. -2x + 8y = 5 8y = 2x + 5 y = 4x + 5. You can see they have the same slope, and this set is parallel. Do the same exact thing for your second problem, and please, show your work, so if you make a mistake, I can correct it! Cheers ~

OpenStudy (anonymous):

1.perpendicular 2.parallel